### v1

# What does really matter in image goal navigation?

Abstract
Image goal navigation requires two different skills: firstly, core navigation skills, including the detection of free space and obstacles, and taking decisions based on an internal representation; and secondly, computing directional information by comparing visual observations to the goal image. Current state-of-the-art methods either rely on dedicated image-matching, or pre-training of computer vision modules on relative pose estimation. In this paper, we study whether this task can be efficiently solved with end-to-end training of full agents with RL, as has been claimed by recent work. A positive answer would have impact beyond Embodied AI and allow training of relative pose estimation from reward for navigation alone. In a large study we investigate the effect of architectural choices like late fusion, channel stacking, space-to-depth projections and cross-attention, and their role in the emergence of relative pose estimators from navigation training. We show that the success of recent methods is influenced up to a certain extent by simulator settings, leading to shortcuts in simulation. However, we also show that these capabilities can be transferred to more realistic setting, up to some extend. We also find evidence for correlations between navigation performance and probed (emerging) relative pose estimation performance, an important sub skill.
1 Introduction
The interplay between perception and action in modern AI-based robotics is a well-studied topic covering contributions from multiple fields, including computer vision, sequential decision making, and optionally some form of mapping and planning. Independent of the specific robotics task at hand, there is an ongoing debate on whether these components should be trained end-to-end, e.g. with RL [8, 75], typically for higher performance, or whether modular techniques could provide higher robustness, e.g. [15, 38].
In the context of navigation, the ImageNav task adds an additional complexity, as perception not only involves the currently observed image, but also the goal image, requiring the comparison of two images to extract directional information. Different strategies have been proposed to address this, and the recent literature seems to provide conflicting reports on their performances and up- and downsides. We aim to shed some light on this with an in-depth analysis.
Figure 1: Image goal navigation requires general navigation skills, but also in particular the extraction of directional information towards the goal. We analyze which architecture design choices influence these capabilities, and to what degree they — and the underlying sub task of relative pose estimation, which we probe with a dedicated head $p$ — can be trained end-to-end from the navigation loss directly, without any pose ground-truth.
Early work addressed the problem in a purely data-driven way by end-to-end training Late Fusion (eg. Siamese) image encoders [77], with mixed results. It was conjectured, that the weak and sparse learning signal from RL training could potentially not be sufficient to train high capacity encoders comparing images. Since the ImageNav task introduces a sub task of relative pose estimation (“where is the goal with respect to my current position?”), there is a strong case for modular methods separating perception and decision taking. For instance, Krantz et al. [38] employ local features and matching (SuperGlue [58]) combined with map and plan strategies for navigation. However, it has since been shown [8] that image correspondences can actually emerge directly in the cross-attention layers of a binocular transformer, provided it is appropriately pre-trained. This strategy, used in the DEBiT agent [8] seems to outperform the explicit matching approaches. As reported in [8], key to this solution is, firstly, pre-training, and secondly, the fusion of observation and goal in such a way that comparisons between local parts (e.g. patches) of the two respective images can be calculated by the architecture, as opposed to comparing embedding vectors.
More recently, some work seems to indicate that ImageNav can be addressed by very low-capacity convolutional encoders (ResNet9) with channel-wise stacking and through RL-training alone, without any pre-training. The FGPrompt model [67] achieves results which are almost on par with DEBiT with a fraction of the model capacity, which raises interesting questions: can directional information be learned by RL losses alone, and by comparably simple architectures? The implications could potentially go beyond image goal navigation and have impact on the wider field of relative pose estimation and localization.
In this paper we disentangle the different mediators and ask fundamental questions on what choices matter in this topic. We find that (i) Certain network architectures better support implicit correspondence computations, that this directly impacts results, with early patch-wise fusion being essential, compared to late fusion. (ii) We show that recent success of frugal architectures using channel stacking is linked to a simulator setting that allows agents to slide along walls. Interestingly, while this setting is known to hamper real-world transfer, to our surprise, some learned capabilities can still be transferred to realistic settings if the transfer includes weights of the perception module. (iii) We also find correlations between navigation and (emergent) relative pose estimation performance, which we probe from representations trained with RL.
2 Related Work
Navigation from visual observations – has been addressed in multiple different fields. Robotics for a long time focused on explicit modeling [13, 43, 45], which is mainly based on mapping and localization [12, 39, 68], explicit planning [37, 62] and low-level control [26, 57]. Accurate sensor and observation models are essential, followed by filtering, dynamical models and optimization techniques.
Modern AI-based solutions are typically trained on large-scale photo-realistic simulators like Habitat [59] or AI2-Thor [36]. Modular agents [17] decompose the problem in sub modules, typically mapping, pose-estimation, planning and local decision taking. On the other end of this spectrum, end-to-end trained models directly map input to actions, with Reinforcement Learning (RL) [31, 48, 76, 69] or Imitation Learning (IL) [23], offline-RL [66], or unsupervised RL [35, 42]. In recent work, end-to-end trained agents are combined with visual foundation models, either general ones like DINOv2 [50] in PoliFormer [76], or binocular ones for image comparisons as in DEBiT [8]. Agent memory is typically implemented as flat recurrent representations [10, 32], occupancy maps [17], semantic maps [16], latent metric maps [5, 30, 51], topological maps [6, 18, 63], scene graphs [65], explicit episodic memory [20, 24, 25, 54], implicit representations [46] or navigability [9]. In this work we investigate agents trained end-to-end with RL and recurrent memory.
Image goal navigation – or “ImageNav”, adds a matching aspect to navigation, as the agent needs to compare the goal image to the observed image. Explicit methods have addressed this with local feature matching [38], or by retrieving features from a topological map [6]. End-to-end trained agents compare images by extracting features with ResNets [1, 77] or ViTs [75], potentially stacking multiple observations over time [64], with ViTs followed by cross-attention [8], or by stacking images channel-wise [67]. Modular approaches have also been proposed [22, 73].
Depending on whether goal images are randomly chosen or correspond to images of semantically meaningful objects, the task can either be supported by semantic features or requires purely geometric image understanding. In this work We focus on the latter, and study how success depends on architectural key design choices. We review these choices in detail in Sec. 3.1 and their implementations in the state of the art in Sec. 3.2.
Relative pose estimation – was tackled for decades with pixel level image matching techniques [28, 49, 60]. Learning-based approaches have also been proposed [34]: to overcome the lack of generalizability of the early methods, self-supervised techniques have been used [41]. More recently, DUSt3R [70] regresses pointmaps of each image expressed in the coordinate system of the first image while MASt3R [40] additionally learns an extra descriptor inspired by standard image matching. Both leverage CroCo as pre-training [71, 72]. Similarly, MicKey [4] regresses pointmaps of each image in its respective coordinate system and learns descriptors to obtain correspondences: this is supervised by relative pose alone with differentiable RANSAC formulations [7, 11]. All these recent methods have led to impressive results for relative pose estimation, even under scenarios with little overlap between input images as in the MapFree-Relocalization benchmark [3]. In this work, we study rel. pose estimation as a sub-task of ImageNav and probe it explicitly from agent representations.
Figure 2: Different architecture choices for binocular encoders learning to compare the observed image $\mathbf{o}_{t}$ with the goal image $\mathbf{g}$: (a) Late Fusion encodes them separately and comparison is done “late” between embedding vectors $\phi_{o}(\mathbf{o}_{t})$ and $\phi_{g}(\mathbf{g})$, making correspondence computations difficult. (b) ChannelCat stacks images over the channel dimension, followed by convolutional encoders $\phi\left([\mathbf{o}_{t},\mathbf{g}]_{\textrm{dim}=1}\right)$. It makes correspondence computations possible in principle if the CNN receptive field is big enough. (c) SpaceToDepth reshapes the patch dimension into the channel dimension. Combined with ChannelCat [67], it could allow correspondence to emerge in each layer directly through conv filters. (d) Binocular ViTs [8] model correspondence directly as cross-attention between patch tokens.
3 Can we learn goal-perception with actions?
In this work we study the ImageNav task in photo-realistic 3D environments, where an agent is given a goal image $\mathbf{g}{\in}\mathbb{R}^{3{\times}H{\times}W}$ and is required to navigate from a starting location to the position shown in the goal. At each timestep $t$ the agent observes an RGB image $\mathbf{o}_{t}{\in}\mathbb{R}^{3{\times}H{\times}W}$, all images are of size $112{\times}112$. The action space is the discrete set $\mathcal{A}$ ={MOVE FORWARD 0.25m, TURN LEFT $10^{\circ}$, TURN RIGHT $10^{\circ}$, and STOP}. An episode is considered successful if the agent is within 1m of the goal position and it calls the STOP action within its 1000 steps budget. We use the Habitat simulator [59] and the Gibson dataset [74].
[…]
$\begin{array}[]{lll}\tilde{\mathbf{g}}_{t}&=\phi(\mathbf{o}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// binocular perception}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{g}}_{t},\zeta(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// state update}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// policy}}\\ \end{array}$ (1)
where $h$ is the function updating the hidden state $\mathbf{h}_{t}$ of a GRU [21] and gating equations have been omitted for brevity; $\phi$ and $\zeta$ are trainable encoders, and $\pi$ is a linear policy. All agent variants are trained from scratch with PPO [61] with a reward function defined as in [71], which re-uses the definition from [19] for PointGoal,
$r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$
where $\lambda{=}0.01$ is a slack cost to encourage efficiency, $K{=}10$, and $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal.
3.1 Architecture design choices
At the heart of the problem is the sub task of learning to compare images $\mathbf{o}_{t}$ and $\mathbf{g}$ to infer directional information related to relative pose estimation. In the classical ImageNav task, goal images can be taken from any position and do not necessarily show a semantically meaningful object. Given the possible absence of semantic cues, which would have allowed to compare images on a global level, we conjecture that successful comparison requires the computation of some form of correspondence between local parts of the respective images. While pose estimation methods based on local feature matching perform this explicitly [55, 58], recent work have shown that this kind of correspondence can emerge implicitly from large-scale training without explicit supervision of matches, e.g. in CroCo [71] for pairwise problems or DUSt3R [70] involving multiple images and addressing SfM tasks.
In the context of robotics and navigation, the question arises whether it is possible to go a step further, and forgo pose-related losses completely, training image comparisons end-to-end together with the navigation agent from RL alone. This has been the promise of works like FGPrompt [67] and OVRL-v2 [75], and we will analyze this in-depth.
Starting from the premise, that correspondence computations are necessary to extract directional information, we conjecture that these computations need to be supported by the underlying network architectures. We will first describe the basic building blocks for image comparison presented in the literature, and also illustrated in Figure 2:
Late fusion Networks
use separate networks $\phi_{o}$ and $\phi_{g}$ to encode observation and goal. The representation fed to the agent is given as $\phi(\mathbf{o}_{t},\mathbf{g})=[\phi_{o}(\mathbf{o}_{t}),\phi_{g}(\mathbf{g})]$, where [.][.][ . ] denotes concatenation, see Figure 2(a). The comparison is thus performed “late” between embeddings $\phi_{o}(\mathbf{o}_{t})$ and $\phi_{g}(\mathbf{g})$, which makes it generally harder to be done on a local image level, unless output representations retain sufficient spatial structure going through encoder layers.
ChannelCat
involves the usage of a single network to encode both, observation and goal, which are channel stacked into one input image, denoted as $\phi\left([\mathbf{o}_{t},\mathbf{g}]_{\textrm{dim}=1}\right)$, where 0 is the batch dimension and 1 the channel dimension — see Figure 2(b).
If encoders are implemented as CNNs, a common choice in robotics for computational complexity reasons, comparisons between local parts of the respective images can be done only if the receptive field of the network is big enough to encompass the full image. Given a sufficiently deep encoder, correspondence can therefore be computed, but not necessarily efficiently, as individual matching results need to be accumulated in the output of different filter kernels.
If encoders are ViTs, each feature patch corresponds to a pair of patches having the same coordinates in both images. Correspondence can be more easily computed if different self-attention heads are learned with query and key projections focusing on one of the two channel groups of the two respective images [^2] As an illustrative example, consider an input image $X$ corresponding to channel stacking of two images $X_{A}$ and $X_{B}$. Then, if for a given attention head the Query projection only uses the channels of image $X_{A}$ and the Key projection only use the channels of image $X_{B}$, then this self-attention on $X$ mimics cross-attention between $X_{A}$ and $X_{B}$..
SpaceToDepth
reshapes image patches into channel values and was introduced in  [56]. FGPrompt [67] uses SpaceToDepth in combination with Channel stacking observation and goal, achieving strong navigation results. We will investigate whether a ResNet in this configuration can compute correspondences across a large spatial dimension through a single convolutional layer, somewhat reminiscent of cross-attention, only with very few parameters — see Figure 2(c).
Cross-attention
is a natural way to compute correspondences between local parts of images [8], as each patch in one image can be naturally linked to one or more patches in the other image through the cross-attention distribution, see Figure 2(d).
3.2 Implementations in prior work
Early works used late fusion approaches based on CNNs [1, 77], but later work switched encoders to ViTs, for instance OVRL-v2 [75]. Cross-attention for ImageNav was introduced in the DEBiT encoder [8], with code made available [^3] https://github.com/naver/debit, and combined with pre-training, first for cross-view completion as in [71], then for relative pose and visibility estimation. The binocular ViT was combined with an additional CNN taking only the observation as input, trained with RL. It currently provides the state-of-the-art results on this task. Interestingly, the authors have shown that pre-training is essential, as the high-capacity binocular transformer cannot be trained from scratch with RL alone, leading to performance close to zero. The RL loss seems to be too weak of a learning signal, not sufficient to drive the encoder to discover correspondence computations. The inability of RL to learn big encoders was somewhat confirmed in the Late Fusion method OVRL-v2 [75], as “switching from the 50.9M parameters ViT-Small to the 179.2M parameters ViT-Base produces another negative result: SR drops -2.4% while SPL minimally increases by +0.7%”.
FGPrompt [67] first introduced channel stacking for ImageNav, with code made available [^4] https://github.com/XinyuSun/FGPrompt. It is heavily based on the previous code of ZSEL [1]. Both codebases use a non-standard ResNet9 architecture that adopts several custom blocks, the most significant one being the SpaceToDepth modification, described earlier. Interestingly, SpaceToDepth was first introduced in TResNet [56] as a general purpose network block for computer vision optimized for memory efficiency, targeting image classification and object detection. Correspondence problems were not explicitly targeted. The ResNet9 encoder in this paper is exactly the same as used in ZSEL and FGPrompt.
Given the good results FGPrompt obtained without any pre-training, 92% SR, vs. DEBiT’s 94% SR depending on pre-training, we raise the question whether it is possible to use the weak learning signal provided by RL to learn binocular perception modules capable of providing directional information. A positive answer would have significant impact also on training vision modules targeting RPE directly. In the experimental section we will dive into this question and benchmark architectures and design choices.
4 Experimental Setup
Unfortunately there are multiple different evaluation protocols in the literature on Embodied AI, and in particular on the ImageNav task. We therefore train all agents in identical conditions and with the same experimental protocol. The agents are trained on the 72 scenes of the Gibson dataset [74] and use the standard Habitat episode definitions (a subset of the literature uses the definitions from Episodic Transformers [47] and another subset uses the definitions from NSNRL [27]). We set the maximum episode length to the default value of 1000 steps and we train for 500M steps, which allows all agent variants to converge easily.
Implementation details – We test perception networks $\phi$ implemented as ResNet, ViT [^5] https://github.com/huggingface/pytorch-image-models and DEBiT architectures, available in their respective public repositories. For the ResNet, we consider a ResNet9 because of strong performance reported in [67], close to those obtained with a ResNet50, but with a fraction of the parameters. We use the Small version of ViT since [75] reports higher navigation performance than Base (when not pre-trained, as here), again with a fraction of the parameters. The function $\zeta$ encodes previous actions into a 32D feature, $h$ is a GRU with 2 layers and hidden dimension 128, followed by a linear Actor-Critic policy $\pi$ — see the appendix for more details.
Validation – again, there does not seem to be a clear protocol for choosing checkpoints from a training run, as some papers seem to take the last and some report the best validation performance over several checkpoints, sometimes up to several thousands. To keep the experimental protocol sound and avoid overfitting evaluation over the choice of checkpoints, while at the same time staying comparable, we use three different splits. We train on the standard Gibson train split (72 scenes) and report on the standard Gibson validation split (14 scenes), as the literature, but we use an independent custom holdout set of scenes to choose the checkpoint. These additional scenes have been taken from the unused set of the full Gibson set.
Model s2d Backbone PT SR(%) SPL(%)
(a) Late Fusion ResNet9 13.8 8.0
(b) [1] Late Fusion ResNet9 12.5 7.6
(c) Late Fusion ViT-Small 12.5 6.7
(d) ChannelCat ResNet9 83.2 43.9
(e) [67] ChannelCat ResNet9 83.6 42.1
(f) ChannelCat ViT-Small 71.1 34.3
(g) Cross-att DEBiT-b 0.0 0.0
(h) [8] Cross-att DEBiT-b 90.5 60.3
Table 1: Comparisons of agents with different binocular visual encoders, trained and validated with Sliding=True. s2d=SpaceToDepth, not applicable to DEBiT and ViT. PT=pre-training for relative pose and visibility estimation.
Metrics – Navigation performance is evaluated by success rate (SR), i.e., fraction of episodes terminated within a distance of ${<}1$m to the goal by the agent calling the STOP action, and SPL [2], i.e., SR weighted by the optimality of the path, $\textit{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$ where $S_{i}$ be a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
Comparing the comparable – An important choice in the Habitat simulator is the binary Sliding switch, which, when switched on, allows the agent in simulation to slide along obstacles when colliding, against the more realistic behavior of stopping when colliding. This parameter is known to have a big impact on sim2real transfer [33]. In our own analysis, we saw that the increased difficulty of Sliding=False goes beyond sim2real transfer and also impacts task difficulty in simulation: when the agent hits a wall, it can in principle continue the episode, but is frequently restricted in its motion, often stuck, and typically needs to turn to recover.
While this setting is switched to True as a default, there is consensus in the field that it should be set to False, as the goal is to decrease the sim2real gap and to evaluate agents in situations which are as close as possible to real physical robots and environments. Looking closer into the available code of ImageNav repositories, it appears most use Sliding=False (DEBiT [8], ZSEL [1], ZSON [44], PIRLNav [53]), with a couple of exceptions using Sliding=True, like FGPrompt [67]. We therefore performed experiments with both settings, and were surprised to see that there is a big influence of this parameter. Given its impact, we color-coded the entries in all tables into three different groups: (i) Sliding=True: the visual encoder $\phi$ has been trained with this setting on; (ii) Sliding=False: it has been trained with this setting off; (iii) Pre-train: the visual encoder $\phi$ has been pre-trained on an auxiliary task, typically relative pose and visibility estimation (RPVE).
Model s2d Backbone PT SR(%) SPL(%)
(a) Late Fusion ResNet9 12.8 7.1
(b) [1] Late Fusion ResNet9 13.2 8.9
(c) Late Fusion ViT-Small 6.9 4.5
(d) ChannelCat ResNet9 44.6 23.4
(e) [67] ChannelCat ResNet9 31.7 18.7
(f) ChannelCat ViT-Small 35.3 16.2
(g) Cross-att DEBiT-b 0.0 0.0
(h) [8] Cross-att DEBiT-b 81.7§ 52.0§
[…]
4.1 Benchmarking architectures
Table 1 presents a comparison of different network architectures trained with Sliding=True. We observe that the ChannelCat architecture obtains excellent performance, in SR almost approaching the levels of DEBiT-b, which, however, benefits from a more complex architecture based on cross-attention, and in particular pre-training on cross-view completion and RPVE. Without pre-training, the DEBiT architecture is not exploitable, as also reported in [8]. Both ResNet9 and ViT Late Fusion architectures underperform, which we explain by the simplicity of comparing observation and goal image late, at the outputs of the encoder $\phi$. Adding SpaceToDepth does not seem to have a significant impact neither on ChannelCat nor on Late Fusion ResNet9: being able to compare local patches across images in a single layer through convolutions does not translate into numerical gains.
[…]
4.2 Transferring capabilities obtained w. “Sliding”
The high impact of the Sliding simulator setting came as a big surprise to us. While its role in sim2real transfer of point goal navigation was established in [33], the common conjecture was that it impacted “raw” navigation performance, i.e., finding free space, avoiding obstacles, short, capabilities which are independent of the extraction of goal directional information. Our experiments, however, suggest that the particular choice of binocular visual encoder $\phi$ is impacted by this hyper-parameter, which in theory should be related to motion only.
Perception Action (%)
Checkpoint $\phi$ $h$ $\zeta$ $\pi$ SR SPL
(a) Load all “false” $f$$\ast$ $f$$\ast$ $f$$\ast$ $f$$\ast$ 31.7 18.7
(b) Load all “true” $t$$\ast$ $t$$\ast$ $t$$\ast$ $t$$\ast$ 54.6 27.5
(c) Load all “true” $t$$\rightarrow$ $t$$\rightarrow$ $t$$\rightarrow$ $t$$\rightarrow$ 65.7 34.1
[…]
This then raises the question, whether (i) the two agents trained in their respective settings have similar capabilities but perform differently purely due to the difference in task difficulty, or (ii) whether, for some reason, training with sliding actually leads to different and perhaps better performing agent capabilities. We tested these hypotheses by performing experiments loading the weights of an agent trained with Sliding=True and finetuning it with Sliding=False. For these experiments we used the ChannelCat + SpaceToDepth agent, for which the difference in SR dropped from 83.6% to 31.7% when sliding was switched off (Tables 1 and 2).
Table 3 shows the results of these experiments. Taking the agent trained with Sliding=True and validating it on False, row (b), makes SR jump to 54.6%, compared to the baseline of 31.7%, trained on False only. This is a surprising result. Although the performance is a far cry of the 83.6% obtained by training and validating with True, it indicates that a certain amount of transfer is possible — some capabilities are learned with sliding enabled. Finetuning this agent for 100M steps on False then provides further gains and yields SR=65.7%, row (c).
To pinpoint the effect, we loaded different parts of the agent trained with Sliding=True: only the perception part (weights of the visual encoder $\phi$), only the action part (GRU $h$, policy $\pi$, encoder $\zeta$ of the previous action) or both. Transferring the action part of the agent, ie. GRU $h$, policy $\pi$, encoder $\zeta$ of the previous action, rows (d) and (e), does not lead to any discernible performance: learning the upstream visual encoder $\phi$ from scratch through pre-loaded downstream weights of the GRU and policy seems a very hard task. However, transferring the perception weights $\phi$ and training the action part from scratch, rows (f) and (g), leads to exploitable results, but significantly below loading the full agent weights. With a SR of 38.5% for the finetuning version, row (g), these are still higher than the baseline of 31.7% trained from scratch on False. In Table 4 we perform similar experiments with DEBiT which, however, uses a pre-trained visual encoder $\phi$ that is kept frozen. We load the rest of the agent trained with Sliding=True and evaluate it with False (b): in this case, no transfer happens and performance drops significantly. Finetuning the agent trained with True on False, (c), improves the Sliding=True policy, but not to the extent of reaching the performance of the agent trained on-distribution, i.e. Sliding=False, arguably because RPVE pre-training of perception is already providing a strong visual encoder.
Perception Action (%)
Checkpoint $\phi$ $h$ $\zeta$ $\pi$ SR SPL
(a) Load all “false” $p$$\ast$ $f$$\ast$ $f$$\ast$ $f$$\ast$ 81.7 52.0
(b) Load all “true” $p$$\ast$ $t$$\ast$ $t$$\ast$ $t$$\ast$ 59.5 35.4
(c) Load all “true” $p$$\ast$ $t$$\rightarrow$ $t$$\rightarrow$ $t$$\rightarrow$ 79.6 46.9
Table 4: OOD behavior (DEBiT agent): $f$: load from agent trained with Sliding=False, $t$: load from agent trained with Sliding=True, $p$: pre-trained with RPVE, $\ast$: frozen, $\rightarrow$: finetune.
We attempt to draw the following lessons from these experiments. The difference between Sliding=True or Sliding=False cannot only be described by the task difficulty alone. The easier task (True) allows to learn additional capabilities, which transfer to the harder task, and which are partially related to perception (since the performance in Table 3(g) $>$ 3(a)), and also related to action / sequential decision taking, since 3(b) $\gg$ 3(a) and 3(b) $\gg$ 3(g). However, we conjecture that the transfer of the knowledge stored in the action component is also highly linked to improving the perception skills, as it does not happen when $\phi$ is pre-trained and frozen (Table 4(a) $\gg$ 4(b), and 4(a) $>$ 4(c)). One possible explanation could be that pure training with False leads to undertraining of, both, action and perception: the policy keeps getting stuck (which we empirically confirmed) and does not learn to cope with the last meters of each episode; this, in turn, leads to undertraining the comparison between the (hardly ever seen) goals and observations.
Model s2d Backbone S %corr.poses %corr.vis.
(Table nr. + row) 1m,10° 2m,20° $\mathbf{{<}0.05}$
Late Fusion 1a ResNet9  7.6 26.0 13.8
Late Fusion 1b ResNet9  9.0 29.6 16.1
ChannelCat 1d ResNet9 11.4 29.3 13.7
ChannelCat 1e ResNet9 18.4 41.6 20.8
Late Fusion 2a ResNet9  7.8 26.8 13.2
Late Fusion 2b ResNet9  8.7 28.5 16.1
ChannelCat 2d ResNet9  9.8 26.9 13.8
ChannelCat 2e ResNet9 12.5 31.9 19.2
ChannelCat 3c ResNet9 $\rightarrow$ 18.2 41.4 21.1
ChannelCat 3d ResNet9 $\rightarrow$  5.8 22.9  6.7
ChannelCat 3e ResNet9 $\rightarrow$  7.2 26.1 11.9
ChannelCat 3g ResNet9 $\rightarrow$ 18.6 41.6 21.0
Cross-att 1h,2h DEBiT-b N/A 92.1 96.8 88.8
Late fusion DEBiT-b N/A 14.8 38.6 19.6
Late fusion DINOv2 N/A 12.9 34.0 22.7
Table 5: Probing relative pose and visibility estimation: representations $\phi(\mathbf{o},\mathbf{g})$ trained with an RL (navigation) loss are frozen, and then we train a probing head. The last block of methods in cursive is not comparable, as the encoders were pre-trained: DEBiT [8] was pre-trained on RPVE losses, DINOv2 [50] with SSL. We report RPVE performance on a hold-out set. s2d=SpaceToDepth, S = Sliding=True. The third block shows agents which have been finetuned ($\rightarrow$) from True to False, cf. Table 3.
4.3 Probing RPVE capabilities
Given that a certain limited capacity for visual reasoning has been learned by the visual encoder $\phi$ when the full agent has been trained for navigation with Sliding=True, we pursue this question further and directly investigate how well the different visual encoders can deal with extracting directional information. Or, motivated differently, given the potential impact on relative pose estimation of training these visual encoders from navigation losses only without pose supervision, we evaluate how well they work for this task. We took the frozen visual encoders $\phi$ of the agents evaluated in Tables 1 and 2 and trained probing heads $p$ on top of them predicting relative pose and visibility (RPVE):
$(\mathbf{t},\mathbf{R},v)=p\Big{(}\phi(\mathbf{o},\mathbf{g})\Big{)}.$ (2)
In the lines of [8], relative pose between the observed image and the goal image is composed of two components, translation $\mathbf{t}{\in}\mathbb{R}^{3}$ and a rotation matrix $\mathbf{R}{\in}\mathbb{R}^{3\times 3}$. Visibility quantifies the amount of overlap between the two images, necessary since in navigation settings the goal image might not even be observed at certain moments. It is defined as the proportion of $16{\times}16$ patches of the goal image $\mathbf{g}$ which are visible in the observed image $\mathbf{o}$.
The probing heads $p$ are adapted to the different architectures, while at the same time providing a comparable capacity of $\sim{3M}$ parameters for each head variant. They collect features at the last spatial representation (feature map for ResNet, patch embeddings for ViT, concatenation of these for Late Fusion representations) and first token-wise linearly project to a lower dimension, before flattening to obtain a global representation which is passed to an MLP with 1024 hidden units. The low dimension of the projection changes according to architectures (as the spatial resolution, i.e., the number of tokens change) to make the parameter counts similar: 64 for ViTs, 192 (resp. 16) for ResNet9 with (resp. without) SpaceToDepth.
Refer to captionDEBiTLFCCCC
Refer to captionDEBiTCC,Load allCC,Load perceptionCCCCLFCC,Load action
(a) Nav evaluated w. Sliding=True         (b) Nav evaluated w. False
Figure 3: Nav vs. Rel-pose: navigation perf. (SR,%) plotted against pose est. probing accuracy (for err ${<}2m,20^{\circ}$) for 4 types of visual encoders $\phi$: trained w. sliding, trained w/o sliding, pre-trained w. RPVE, trained w. sliding, finetuned w/o; (LF = Late Fusion, CC=ChannelCat). The dashed line relates the finetuned models to the same model trained w/o sliding.
Figure 4: Analysis of navigation behavior: Sankey plots show the distribution of success/failure codes over 944 test episodes for different models, and their “flow” between certain pairs of models. For instance, the strength of the connection between “Time out” (left) and “success” (right) indicates how many episodes toggled from one to the other when switching from the left to the right model.
We generated a probing dataset by combining the 3D scene datasets Gibson [74], MP3D [14], and HM3D [52], following their standard train/val scene splits, as done in  [8]. We sample pairs of points randomly in the scene, calculate the shortest path between them, then extract 10 intermediate poses on the path. These 10 poses are then each combined with the end point to form pairs of observation $\mathbf{o}$ and goal $\mathbf{g}$. GT for relative pose is directly available from the simulator, and visibility can be estimated from the two point clouds generated from the respective images. We generated around 68M image pairs total.
We train the probing heads $p$ with a loss combining all three components, translation, rotation and visibility, where pose supervision is switched off when visibility is low,
$\mathcal{L}_{p}=\sum\mathop{}_{\mkern-5.0mui}\Bigl{[}|v_{i}{-}v_{i}^{*}|+\mathbf{1}_{v^{*}_{i}{>}\tau}\bigl{\{}|\mathbf{t}_{i}{-}\mathbf{t}_{i}^{*}|+|\mathbf{R}_{i}{-}\mathbf{R}_{i}^{*}|\bigr{\}}\Bigr{]},$ (3)
where $i$ indexes image pairs over the probing dataset, $\mathbf{t}_{i}^{*},\mathbf{R}_{i}^{*},v_{i}^{*}$ denote ground truth values, $\mathbf{1_{.}}$ is the binary indicator function, |.||.|| . | denotes the $L_{1}$ loss and $\tau$ is a threshold.
Metrics – relative pose is evaluated over the pairs with visibility over $\tau$ in the percentage of correct poses for given thresholds on distance and angle, e.g. 1 meter and 10°. Visibility is evaluated over all pairs by its accuracy at $\pm$0.05, i.e., the percentage of prediction within a 0.05 margin of the ground-truth value.
Results – Table 5 gives results comparing RPVE performance on a validation set for all tested visual encoders, and Figure 3 provides a scatter plot relating it to navigation performance (taken from Tables 1, 2, 3). While the ChannelCat architecture obtains high navigation performance in the Sliding=True setting, 3(a), its pose estimation performance remains limited. Late Fusion models generally provide low performance both for navigation and pose estimation. In the Sliding=False setting, 3(a), the navigation performance significantly drops for the ChannelCat models, as already discussed in Section 4.1, and also somewhat for pose estimation. Transferring the models trained with True on the False setting provides gains not only in navigation performance, as already discussed, but also on pose estimation — only when at least the weights of the visual encoder $\phi$ are transferred, but in particular when the whole agent is transferred and finetuned. This further corroborates our conjecture, that the perception model is undertrained when sliding is disabled. Finally, the best performance is obtained by the DEBiT model based on ViTs and cross-attention, both in navigation and pose probing. However, the pose estimation performance is not comparable for obvious reasons — the model has been pre-trained and frozen for this task. Its performance is given for information.
In Table 5 we also provide pose probing results on two additional models, where we explore whether the pose estimation performance of late fusion can be increased by using two different pre-trained and frozen encoders, namely DINOv2 and the ViT encoder part of DEBiT, without the cross-attention layers combining them [^6] DEBiT first encodes each image with a ViT and then passes both embeddings to a decoder using cross-attention, cf. [8].. They both underperform, further corroborating the importance of comparing visual representations early, allowing to compute correspondences between local parts of images.
4.4 Analysis of navigation behavior
In Figure 4 we provide Sankey plots showing the differences in navigation behavior as a distribution of success/error types over 994 test episodes evaluated in environments with Sliding=False, and how this behavior varies across different models and different training settings. Here we distinguish between unsuccessful episodes due to Time out, when the agent does not complete an episode in the 1000 steps budget, and Early stop, when the agent terminates the episode but is further away than the 1m success threshold.
Left, the ChannelCat agent trained on Sliding=False, Table 3(a), stops too early in most of the episodes, confirming that its components are not properly trained, probably due to the difficulty of the task. Next, finetuning on False the ChannelCat agent trained with True, Table 3(c), allows to obtain an agent with good navigation capabilities. It improves the ChannelCat agent trained with True, Table 3(b), in the middle in Figure 4, whose main failure mode is time out. We observe that most of these time outs occur far from the goal and are due to the agent getting stuck, likely because it learned to slide along obstacles, which is not possible with False. On the 4th column, DEBiT trained with True, Table 4(b), exhibits a similar behavior, while DEBiT trained with False, Table 4(a), on the last column, can effectively solve the complex navigation task, likely thanks to its pretrained, high capacity, perception module.
5 Discussion and conclusion
In this paper we have studied the connections between image goal navigation and relative pose estimation. We have shown that the success of recent late fusion architectures based on channel stacking can be linked to a simulator setting allowing the agent to slide along walls. While agents trained with sliding enabled are known to transfer badly to real environments, where this is not possible, our analysis has shown that capabilities learned with this sliding setting can actually be partially transferred to more realistic environment settings. While sliding is a property of agent dynamics and not inherently linked to perception, we have shown that it significantly impacts training of the visual encoders of end-to-end trained agents: transferring capabilities to the realistic setting only seems to be successful if it also involves transferring the visual encoder weights. We conjecture that training with the realistic setting (sliding disabled) leads to undertraining both the perception module, as the goal is seen more rarely, as well as the rest of the agent, which has a difficulty to cover later portions of the episodes.
We also compared and analyzed different widely-used architectures: Late Fusion, ChannelCat, SpaceToDepth projections and Cross-attention. We conjecture that successful architectures require support for early fusion of representations, allowing to compute correspondences between local parts of images, and we found evidence for this in our experiments: Late Fusion generally underperforms. As expected, our experiments also found a correlation between navigation performance and (emerging) relative pose estimation performance, which we probed with explicit heads. Finally, we argue that, up to now and up to our knowledge, there does not seem to exist a simple solution to learn image goal navigation successfully end-to-end with simple relatively low-capacity architectures and without pre-training, which we judge to remain essential for goal oriented navigation.
Appendix A Implementation details
ResNet backbone – The ResNet implementation is borrowed from [56, 1, 67] and the code is publicly available  [^7] https://github.com/XinyuSun/FGPrompt. The default network is implemented as a custom ResNet9 with SpaceToDepth blocks [56] that rearrange the input of each 2D convolution layer such that each of its spatial dimensions is divided by a factor 4, while the channel dimension is 16 times larger. We consider a ResNet9 because of the performance reported in [67], in line with those of a ResNet50, but with a fraction of the parameters. In the ResNet version without SpaceToDepth, these reshaping blocks are simply skipped. As customary, the 2D CNN output is flatten and passed to a linear layer to generate a 128D feature vector.
The Late Fusion ResNet backbone uses two such networks and concatenates the output vectors before feeding them to the GRU, while in the ChannelCat version the first 2D convolution layer is modified to receive as input a 6-channel image generated by channel-stacking observation and goal.
ViT backbone – We use the small version of a popular ViT implementation [^8] https://github.com/huggingface/pytorch-image-models, augmented with the convolutional compression layer proposed in OVRL-v2 [75] (code available in the paper) to compress the ViT output to a 2058D feature. As discussed in [75], without pre-training (as is the case here), ViT-Small performs better than Base despite having a fraction of the parameters, we thus use ViT-Small for all our experiments. We adopt this layer since Yadav et al. [75] report significant improvements in navigation performance using it, over more common compression options like average pooling or [CLS] token.
As for the ResNet backbone, the Late Fusion network is implemented with two ViTs whose outputs are concatenated, while the ChannelCat version has a modified input 2D convolution (first standard processing layer of the ViT) such that it can receive a 6-channel image composed by channel-stacking observation and goal.
DEBiT backbone – In the paper we use the DEBiT-b version of the architecture, the variant of DEBiT with the best performance-size trade-off and publicly available with its pre-trained binocular encoder  [^9] https://github.com/naver/debit. The binocular encoder is kept frozen and is based on the Geometric Foundation Model CroCo [71], finetuned to estimate relative pose and visibility as described in [8] and the main text. The network is implemented with two Siamese encoders, applied to observation and goal images, and a decoder which combines the output of these two encoders. Both encoders and decoders are implemented with a ViT architecture with self-attention layers, with the decoder also featuring cross-attention. The output of the decoder is further compressed by a fully connected layer (also kept frozen, from the DEBiT model) that projects the flattened output of the decoder into a 3136D feature vector.
The DEBiT model also features an observation encoder, needed since the binocular encoder is frozen and can thus only perform observation-goal comparison, without fulfilling other tasks needed for navigation such as, for example, free space estimation. It is implemented with a half-width ResNet18 [29] that generates a 512D output feature.
Model+backb. GRU architecture Sliding SR(%) SPL(%)
DEBiT-b 2 layers, $|\mathbf{h}_{t}|=128$ 81.7 52.0
DEBiT-b 2 layers, $|\mathbf{h}_{t}|=128$ 90.5 60.3
DEBiT-b 1 layer, $|\mathbf{h}_{t}|=512$ 83.7 54.8
DEBiT-b 1 layer, $|\mathbf{h}_{t}|=512$ 90.8 60.5
Table 6: Comparisons of DEBiT [8] in two settings: our standard 2-layer GRU test architecture used in the paper (Tables 1,2 and 3), and the architecture originally reported in [8] (1-layer GRU).
Appendix B Additional experiments
Impact of GRU structure – For fairness of comparison, we used the same GRU architecture in all our experiments, with 2 layers and hidden state of size 128. In Table 6 we show the results for training DEBiT-b with the same GRU architecture which reported in [8], with one layer and hidden dimension 512. The experiment excludes any dependency on this choice: the results are very similar in spite of the different number of layers and the change in sizes of the recurrent memory $|\mathbf{h}_{t}|$.